ITES2001 Fourth Italian-Spanish conference on General Topology and its applications Bressanone, 27-30 June 2001

Generating quasi-metrics from norms on semilinear spaces

S. Romaguera, E. A. Sánchez-Pérez, O. Valero Departamento de Matemática Aplicada Universidad Politécnica de Valencia, Spain

We present a method to generate quasi-metrics from norms defined on semilinear spaces. The Sorgenfrey quasi-metric on [0,+∞) and a new complexity distance are obtained as particular cases of such a method.